• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name V. Sulkosky

    Log entry time 18:56:11 on January26,2005

    Entry number 140104

    keyword=Loop 1 Temperature Sensor!

    17:45 While I was moving the target to 15 cm LD2, the Loop 1
    temperature sensor apparently went bad. Currently it is reading
    about 120 K. I paged JP, and he changed the heater control to manual
    and set the HPH to adjust the remaining temperature sensors to 22 K.

    I also paged Chris Keith to change the temperature sensor that the
    PID feedback uses to cernox 2. Once this was done, I placed the
    heater control back to PID. The target temperature is now stable at
    22 K. However the temperature on the TV screen and the main GUI
    still read from the main sensor, which is 120 K. From the charts of
    Loop 1 in Figure 1, target operators should look at the second (green)
    , fourth (blue), and fifth (purple) temperture sensors.

    Target operators need to pay special attention to the Loop 1
    temperature, since some of the read backs are incorrect!

    FIGURE 1